OpenStudy (anonymous):

Find the slope of the line described by 8x = 48 + 3y.

OpenStudy (anonymous):

slope= 8/3

OpenStudy (anonymous):

The easiest way to determine the slope of a linear equation is by rearranging it so that it is in slope-intercept form:\[y = mx +b\] For this equation, it would be like this\[48+3y=8x\]Subtract 48 from both sides\[3y = 8x-48\]Then divide both sides by 3:\[y = \frac{8}{3}x-16\]Now that it is in slope-intercept form, you can know that the coefficient of x (that is, m) is the slope. In this case, 8/3

OpenStudy (anonymous):

^ thanks for typing out that whole explanation... if you're taking a test, you need to know that formula.